OpenStudy (anonymous):

cos(3π/8) Without a calculator. has to be exact…

OpenStudy (watchmath):

First we need to know \(\cos(\pi/8)\). Let \(x=\pi/8\). \(\cos(2x)=2\cos^2(x)-1\) On the other hand \(\cos(2x)=\cos(\pi/4)=\frac{1}{2}\sqrt{2}\) It follows that \(2\cos^2(x)-1=\frac{1}{2}\sqrt{2}\) \(\cos x=\frac{1}{2}\sqrt{1+\frac{1}{2}\sqrt{2}}\). Then compute \(cos(3\pi/8)\) by using \(\cos(3\pi/8)=\cos((\pi/8)+(\pi/4))=\cos(\pi/8)\cos(\pi/4)-\sin(\pi/8)\sin(\pi/4)\).